OpenStudy (anonymous):

let log(D^2/C^3*A) when logA=3 logB=2 logD=5 a)0.549 b)4.44 c)1 d)19

OpenStudy (anonymous):

\[\begin{align*}\log\left(\frac{D^2}{C^3A}\right)&=\log D^2-\log\left(C^3A\right)\\ &=2\log D-\left(\log C^3+\log A\right)\\ &=2\log D-3\log C-\log A\\ &=\cdots \end{align*}\]

OpenStudy (anonymous):

i did that but i got none of the four answers that they give you. please explain more

OpenStudy (anonymous):

Do you know what \(\log C\) is?

OpenStudy (anonymous):

logC is 2

OpenStudy (anonymous):

Ah, so that was a typo.. \[\begin{align*}\log\left(\frac{D^2}{C^3A}\right)&=\log D^2-\log\left(C^3A\right)\\ &=2\log D-\left(\log C^3+\log A\right)\\ &=2\log D-3\log C-\log A\\ &=2(5)-3(2)-3\\ &=10-6-3\\ &=\cdots \end{align*}\]

OpenStudy (anonymous):

yeah it was. thanks for the help

OpenStudy (anonymous):

you're welcome